What is the x-intercept of this line? Y+4=-4(x-2)
Answer by Alan3354(69443)

What is the x-intercept of this line? Y+4=-4(x-2)
---------
At the x-int y = 0
Y+4=-4(x-2)
0+4=-4(x-2)
4=-4(x-2)
1 = -(x-2)
-1 = (x-2)
x = 1
or (1,0)